Answer the next 3 questions based on the following information: Assume that3-month Treasury bill are yielding 1.4%, 10-year Treasury bonds are yielding 2.8%, an Aaa-rated 10-year corporate bond is 5%, and real risk-free rate is 1%. 18) What is the inflation risk premium for 3-month Treasury b ill? (Hint: the inte rest rate of 3-month Treasury bill is a proxy for risk-free rate) A) 1.8% B) 1.4% C) 1% D) 0.4% 19) What is the maturity risk premium for 10-year Treasury bond? A) 1.8% B) 1.4% C) 1% D) 2.2% 20) If the liquidity risk premium on this Aaa-rated corporate bond is 0.4%. What is the default risk premium for this corporate bond?A) 1.8% B) 1.4% C) 1% D) 2.2%
Answer #1

18.

Inflation risk premium for 3-month T bill = T bill yield - Rf

Inflation risk premium for 3-month T bill = 1.4 - 1 = 0.4%

So, Option D is correct.

19.

Maturity Risk Premium for 10 year Bond = Bond Yield - Rf - Inflation Risk Premium

Maturity Risk Premium for 10 year Bond = 2.8 - 1 - 0.4 = 1.4%

So, Option B is correct.

20.

Default risk premium of 10 year corporate bond = Bond Yield - Rf - Inflation Risk Premium - Liquidity Risk Premium

Default risk premium of 10 year corporate bond = 5 - 1 - 0.4 - 1.4 = 2.2%

So, Option D is correct.